# Multifrequency Radio Observations of the Magnetar Swift J1818.0–1607

### v1

@@ Abstract
We report on Green Bank Telescope observations of the radio magnetar Swift J1818.0–1607 between 820 MHz and 35 GHz, taken from six to nine months after its 2020 March outburst. We obtained multi-hour observations at six frequencies, recording polarimetric, spectral, and single-pulse information. The spectrum peaks at a frequency of $5.4\pm 0.6$ GHz, making Swift J1818.0–1607 one of many radio magnetars which exhibit a gigahertz-peaked spectrum (GPS). The radio flux decays steeply above the peak frequency, with in-band spectral indices $\alpha<-2.3$ above 9 GHz. The emission is highly ($>50\%$) linearly polarized, with a lower degree ($<30\%$) of circular polarization which can change handedness between single pulses. Across the frequency range of our observations, the time-integrated radio profiles share a common shape: a narrow “pulsar-like” central component flanked by “magnetar-like” components comprised of bright, spiky subpulses. The outer profile components exhibit larger degrees of flux modulation and flatter spectral indices when compared to the central pulse component.
@@ 1 Introduction
Pulsed radio emission has only been detected from six magnetars [^1] https://www.physics.mcgill.ca/~pulsar/magnetar/main.html including the most recent discovery, Swift J1818.0–1607. In the radio band, magnetar emission tends to be highly linearly polarized, with a flat polarization position angle (PPA) sweep across the center of the profile (Kramer et al., 2007). The emission is composed largely of short, spiky ‘sub-pulses’ falling within a larger pulse envelope (Levin et al., 2012). Magnetar emission can possess flat or even inverted spectral indices over large intervals of the radio spectrum, as in the case of XTE J1810–197 (Eie et al. 2021; Torne et al. 2022) and the Galactic center magnetar SGR J1745–2900 (Torne et al. 2015, 2017). This is at odds with the majority of rotation-powered pulsars, whose flux follows power-law spectra with an average spectral index of --1.6 (Jankowski et al., 2018).
@@
Like other radio magnetars, J1818 exhibits changes in the shape and number of components in its integrated profile, high degrees of linear polarization, a variable spin-down rate, and flux variability across a range of timescales (Lower et al. 2020; Huang et al. 2021; Rajwade et al. 2022). Its X-ray properties post-outburst were also reminiscent of other magnetar outbursts (Blumer & Safi-Harb 2020; Hu et al. 2020). More recently, the radio flux of J1818 increased from 2022 December to 2023 March, but subsequently decayed, and no detections have been published since then, suggesting J1818 may have entered a quiescent period after its outburst (Ding et al., 2024).
@@ 2 Observations and Data Reduction
Instances of terrestrial radio frequency interference (RFI) were identified and excised using the paz and pazi routines. Frequency channels contaminated with continuous, narrowband RFI throughout the entire observation were excised, as well as individual subintegrations dominated by impulsive, broadband RFI signals. After RFI mitigation, each observation was folded and split into 16 subintegrations, and averaged in frequency to create 16 pulse profiles. Using presto, we generated times of arrival (TOAs) by cross-correlating the total integrated profile with the profile from each sub-integration. The TOAs were passed to tempo2333https://www.atnf.csiro.au/research/pulsar/tempo2/ (Hobbs et al., 2006), where we created a basic timing model using the DE405 ephemeris and the known sky coordinates, fitting only for the spin period.
@@ 3.1 Time-Averaged Pulse Profiles
Figure 1 displays the time- and frequency-integrated pulse profile of the magnetar at each observing frequency. Above a frequency of 1 GHz, the integrated profiles are made up of three distinct components: a precursor (I), narrow main component (II), and wider post-cursor (III). Due to our lack of a phase-connected timing solution, the profiles are all aligned so that the peak of Component II occurs at a rotational phase of 0.25; the 820-MHz profile is aligned so that its peak occurs at the same phase.
@@
We observe mostly flat PA swings across the phase range of the integrated pulse profiles, especially at 820 MHz, in line with the early low-frequency detections of J1818 (Champion et al. 2020; Lower et al. 2020). The PA swing of the 22 GHz observation is much steeper, and inverted with respect to the other observations, but still retains the characteristic S-shaped sweep if a small offset is added to the position angles in Figure 2. Similar changes in the polarization properties of J1818 were seen at 2.6 GHz by Lower et al. (2021), in which case the PA exhibited a steeper, reversed PA sweep on a single epoch of observation before returning to a positive slope.
@@ 3.2 Flux Density and Spectral Indices
Figure 3: The flux measurements in each subband and their corresponding errors are shown in black. The best-fit power law is overlaid in red, with the associated spectral index and 1-$\sigma$ error at the bottom of each panel. The 6-GHz data are fit with a combined power-law and free-free absorption spectral model (see text). No spectral index was obtained when fitting the 35-GHz data.
Figure 2 shows the time-integrated frequency/phase plots, spectral fits, polarization profiles, and linear polarization position angles for each observation in which the magnetar was detected. We see no evidence of astrophysical scintillation in our dataset. Below 22 GHz, the scintillation bandwidth predicted by the NE2001 model is smaller than the bandwidth of a single frequency channel. The scintillation bandwidth predicted by the NE2001 model is 2.4 MHz in the 22-GHz observing band, and 19 MHz in the 35-GHz observing band. Therefore, even though just resolvable, it would be difficult to detect scintillation with sufficient signal-to-noise at these frequencies.
@@
The system temperature $\mathrm{T_{sys}}=\mathrm{T_{rec}}+\mathrm{T_{sky}}$, where the receiver temperature $\mathrm{T_{rec}}=23$ K (“Observing with the GBT”, version 4.0) [^6] https://greenbankobservatory.org/portal/gbt/observing/ and the sky temperature $\mathrm{T_{sky}}$ was calculated at the central frequency of each subband using the pyGDSM Python interface [^7] https://github.com/telegraphic/pygdsm for the Zheng et al. (2017) global diffuse sky model. The gain of the GBT is 2 K/Jy (“The Performance of the GBT”) [^8] https://www.gb.nrao.edu/GBT/Performance/PlaningObservations.htm, and the on-source time is taken from Table 1. The bandwidth ($\Delta\nu$) and off-pulse standard deviation ($\sigma$) are calculated separately for each subband. We add in quadrature a 10% uncertainty on each flux measurement due to the systematic errors involved in using the radiometer equation, due to uncertainties and time dependence in the gain and system temperature.
Flux measurements of the calibrated 22-GHz observation yielded mean fluxes on the order of 50 $\mu$Jy, at odds with the high S/N of its integrated profile and the mean flux of J1818 at all other frequencies. We used the radiometer equation as a validity check, assuming a system temperature of 40 K and using the same method described above to estimate the fluxes and their corresponding errors. The radiometer equation returns mean fluxes in the range of 0.1 to 1 mJy, and given that these measurements made more physical sense, we determine that there was an error in the flux calibration, and use the radiometer fluxes from the 22-GHz observation for our analyses.
@@
The mean flux of J1818 at each frequency is shown in the top panel of Figure 4. The light gray bars denote the frequency ranges of the 9- and 22- GHz datasets, to distinguish them from the other observations. We note that these observations are spread out over the course of roughly 100 days, and given the known flux variability of magnetars over time, these results cannot be used to study the instantaneous broadband spectrum of J1818 nor the time variability of its spectrum.
@@ 3.3 Phase-resolved Spectroscopy
In order to better understand the emission properties of the individual profile components, we also measured the time-averaged spectral index of each component. We defined the on-pulse region of each individual component using the highlighted phase ranges in Figure 1, and calculated mean on-pulse fluxes for each component by using only these relevant parts of the profile. We used these on-pulse fluxes to calculate the spectral indices of the individual profile components at each frequency. The mean fluxes of each component are shown in the bottom panel of Figure 4, and the resulting fitted spectral indices are provided in Table 2. Overall, we find that the spectral indices of component II are steeper (more negative) than those of I and III at most frequencies.
@@ 4.3 Polarization
Figure 12: The distribution of linear polarization fraction (L/I) and circular polarization fraction (C/I) for each single pulse at 820 MHz. 1-$\sigma$ error bars on each point are denoted in light gray.
Figure 13: The distribution of linear polarization fraction (L/I) and circular polarization fraction (C/I) for each single pulse at 9 GHz. 1-$\sigma$ error bars on each point are denoted in light gray.
@@
We examined the degree of linear and circular polarization within each single pulse by measuring the polarized flux and comparing it with the flux of the total-intensity pulse. The errors on the linearly and circularly polarized fluxes are calculated as the off-pulse RMS of the respective polarized pulse profile (i.e. the RMS of the linearly polarized profile is the error on the linearly polarized flux). To examine the distribution of polarization fractions, we first excised pulses from our sample which had polarized fluxes greater in magnitude than the total-intensity flux and pulses with flux measurement errors larger than the flux measurements themselves.
The resulting distributions of single pulse polarization fractions (linear, L/I, and circular, C/I) at 820 MHz and 9 GHz are respectively shown in Figures 12 and 13. In comparison to the polarization fractions of the time-integrated profiles given in Table 2, the pulses at 820 MHz have a lower average L/I and higher average C/I. At 9 GHz, the L/I distribution is wider and there are more highly linearly polarized pulses, but the C/I distributions are similar at both frequencies. In both cases, there is a slight positive correlation between L/I and $|C/I|$.
At both frequencies, a portion of the single pulses have the sign of the circular polarization reversed. Pulse-to-pulse changes in the handedness of the circular polarization have been observed before in radio magnetars alongside structures in the PA which deviate significantly from the RVM (Kramer et al., 2007; Dai et al., 2019). Similar behavior was also observed in a long-period radio transient with similar emission properties to FRBs and magnetars (Men et al., 2025). We show some of the 9-GHz single pulses and their PA swings in Figure 14.
@@ 4.4 Searches for Periodic Substructure
For each of these pulses and their associated quasi-periodicities, we also wanted to measure a statistical significance for each quasi-periodicity. We took each pulse with an identified quasi-periodicity and randomly scrambled the order of its phase bins 1000 times. We took the ACF of each scrambled pulse and identified the maximum value outside of the first and last bin (e.g. the quasi-periodicity of the simulated pulse), and summed the values of the ACF at the maximum value and its harmonics. The simulated ACF sums form a null distribution, and the mean and standard deviation of this distribution can be compared to the summed ACF power of the real quasi-periodicity to obtain a significance ($\sigma$) for the detected quasi-periodicity.
Of the pulses selected for autocorrelation analysis, 118 pulses at 6 GHz and 24 pulses at 9 GHz had detected quasi-periodicities with significances above 4 $\sigma$. The distribution of quasi-periodicities at both frequencies is shown in Figure 16. At 6 GHz, the geometric mean of the detected quasi-periodicities is $4.9\pm 1.4$ ms; at 9 GHz, the mean is $7.0\pm 1.5$ ms, where the 1-$\sigma$ error bars are equal to the geometric standard deviation. The power-law scaling relation derived by Kramer et al. (2024) predicts a mean quasi-periodicity of $1.27\pm 0.08$ ms, while the distribution of their detected quasi-periodicities had a geometric mean of $2.37\pm 0.04$ ms.
The quasi-periodicities detected at 6 GHz by Kramer et al. (2024) were all shorter than 5 ms, at odds with the larger range of quasi-periodicities we observe at both frequencies. We note that at 9 GHz, many of the pulses with longer quasi-periodicities consisted of only two components separated by 5–15 ms, which would also create peaks in the ACF at a lag equal to their separation. It is likely that the combination of sub-pulse structure and the separation of discrete pulse components led to a higher average quasi-periodicity than predicted by the scaling relation. Quasi-periodic substructure is expected to depend only on the rotational period of the neutron star, thus we do not expect that the detected quasi-periodicities will have a frequency dependence.
@@
Figure 16: The distribution of detected quasi-periodicities in the single pulses of J1818 at 6 GHz (top) and 9 GHz (bottom).

### v2

@@ Abstract
We report on Green Bank Telescope observations of the radio magnetar Swift J1818.0–1607 between 820 MHz and 35 GHz, taken from six to nine months after its 2020 March outburst. We obtained multi-hour observations at six frequencies, recording polarimetric, spectral, and single-pulse information. The spectrum peaks at a frequency of $5.4\pm 0.6$ GHz, making Swift J1818.0–1607 one of many radio magnetars which exhibit a gigahertz-peaked spectrum (GPS). The radio flux decays steeply above the peak frequency, with in-band spectral indices $\alpha<-2.3$ above 9 GHz. The emission is highly ($>50\%$) linearly polarized, with a lower degree ($<30\%$) of circular polarization which can change handedness between single pulses. Across the frequency range of our observations, the time-integrated radio profiles share a common shape: a narrow “pulsar-like” central component flanked by “magnetar-like” components comprised of bright, spiky subpulses. The outer profile components exhibit larger degrees of flux modulation when compared to the central pulse component.
@@ 1 Introduction
Pulsed radio emission has only been detected from six magnetars [^1] https://www.physics.mcgill.ca/~pulsar/magnetar/main.html including the most recent discovery, Swift J1818.0–1607. In the radio band, magnetar emission tends to be highly linearly polarized, with a smaller, but nonzero, degree of circular polarization (Kramer et al. 2007; Camilo et al. 2008; Lower et al. 2024). The polarization position angle (PPA) of the emission often follows the S-shaped curve predicted by the rotating vector model (RVM) (Radhakrishnan & Cooke, 1969), though many deviations from this model have been observed as well as changes in the sign of the slope of the PPA (Dai et al. 2019; Lower et al. 2021, 2024). The emission is composed largely of short, spiky ‘sub-pulses’ falling within a larger pulse envelope (Levin et al., 2012). Magnetar emission can possess flat or even inverted spectral indices over large intervals of the radio spectrum, as in the case of XTE J1810–197 (Eie et al. 2021; Torne et al. 2022) and the Galactic center magnetar SGR J1745–2900 (Torne et al. 2015, 2017). This is at odds with the majority of rotation-powered pulsars, whose flux follows power-law spectra with an average spectral index of --1.6 (Jankowski et al., 2018).
@@
Like other radio magnetars, J1818 exhibits changes in the shape and number of components in its integrated profile, high degrees of linear polarization, a variable spin-down rate, and flux variability across a range of timescales (Lower et al. 2020; Huang et al. 2021; Rajwade et al. 2022; Fisher et al. 2024). Its X-ray properties post-outburst were also reminiscent of other magnetar outbursts (Blumer & Safi-Harb 2020; Hu et al. 2020). More recently, the radio flux of J1818 increased from 2022 December to 2023 March, but subsequently decayed, and no detections have been published since then, suggesting J1818 may have entered a quiescent period after its outburst (Ding et al., 2024).
@@ 2 Observations and Data Reduction
Instances of terrestrial radio frequency interference (RFI) were identified and excised using the paz and pazi routines from psrchive (van Straten et al., 2011). Frequency channels contaminated with continuous, narrowband RFI throughout the entire observation were excised, as well as individual subintegrations dominated by impulsive, broadband RFI signals. After RFI mitigation, each observation was folded and split into 16 subintegrations, and averaged in frequency to create 16 pulse profiles. Using presto, we generated times of arrival (TOAs) by cross-correlating the total integrated profile with the profile from each sub-integration. The TOAs were passed to tempo2333https://www.atnf.csiro.au/research/pulsar/tempo2/ (Hobbs et al., 2006), where we created a basic timing model using the DE405 ephemeris and the known sky coordinates, fitting only for the spin period.
@@ 3.1 Time-Averaged Pulse Profiles
Figure 1 displays the time- and frequency-integrated pulse profile of the magnetar at each observing frequency. Above a frequency of 1 GHz, the integrated profiles are made up of three distinct components: a leading component (I), narrow main component (II), and wider trailing component (III). Due to our lack of a phase-connected timing solution, the profiles are all aligned so that the peak of Component II occurs at a rotational phase of 0.25; the 820-MHz profile is aligned so that its peak occurs at the same phase.
@@
The PA swing at 820 MHz is flatter than at higher frequencies, while still retaining the expected S-shaped curve. The relatively flat slope is likely caused by interstellar scattering (Kramer & Johnston, 2008), and agrees with earlier low-frequency detections of J1818 (Champion et al. 2020; Lower et al. 2020). The PA swing of the 22 GHz observation is much steeper, and inverted with respect to the other observations, but still retains the characteristic S-shaped sweep. Similar changes in the polarization properties of J1818 were seen at 2.6 GHz by Lower et al. (2021), in which case the PA exhibited a steeper, reversed PA sweep on a single epoch of observation before returning to a positive slope.
@@ 3.2 Flux Density and Spectral Indices
Figure 3: The flux measurements in each subband and their corresponding errors are shown in black. All of the y-axes are in log scale. The best-fit power law is overlaid in red, with the associated spectral index and 1-$\sigma$ error at the bottom of each panel. The 6-GHz data are fit with a combined power-law and free-free absorption spectral model (see text). No spectral index was obtained when fitting the 35-GHz data. The reduced chi-squared values associated with each fit are given in the top right corner of each sub-panel.
Figure 2 shows the time-integrated frequency/phase plots, polarization profiles, and linear polarization position angles for each observation in which the magnetar was detected. We see no evidence of astrophysical scintillation in our dataset. Below 22 GHz, the scintillation bandwidth predicted by the NE2001 model is smaller than the bandwidth of a single frequency channel. The scintillation bandwidth predicted by the NE2001 model is 2.4 MHz in the 22-GHz observing band, and 19 MHz in the 35-GHz observing band. Therefore, even though just resolvable, it would be difficult to detect scintillation with sufficient signal-to-noise at these frequencies.
@@
The system temperature $\mathrm{T_{sys}}=\mathrm{T_{rec}}+\mathrm{T_{sky}}$, where the receiver temperature $\mathrm{T_{rec}}=23$ K (“Observing with the GBT”, version 4.0) [^6] https://greenbankobservatory.org/portal/gbt/observing/ and the sky temperature $\mathrm{T_{sky}}$ was calculated at the central frequency of each subband using the pyGDSM Python interface [^7] https://github.com/telegraphic/pygdsm for the Zheng et al. (2017) global diffuse sky model. The gain of the GBT, $G$, is 2 K/Jy (“The Performance of the GBT”) [^8] https://www.gb.nrao.edu/GBT/Performance/PlaningObservations.htm, and the on-source time, $t_{obs}$, is taken from Table 1. The bandwidth ($\Delta\nu$) and off-pulse standard deviation ($\sigma$) are calculated separately for each subband. The number of recorded polarizations, $n_{p}$, is two for the full Stokes observations. We add in quadrature a 10% uncertainty on each flux measurement due to the systematic errors involved in using the radiometer equation, due to uncertainties and time dependence in the gain and system temperature.
Flux measurements of the calibrated 22-GHz observation yielded mean fluxes on the order of 50 $\mu$Jy, at odds with the high S/N of its integrated profile and the mean flux of J1818 at all other frequencies. We used the radiometer equation as a validity check, assuming a system temperature of 40 K and using the same method described above to estimate the fluxes and their corresponding errors. The radiometer equation returns mean fluxes in the range of 0.5 to 1.7 mJy, and given that these measurements made more physical sense, we determine that there was an error in the flux calibration, and use the radiometer fluxes from the 22-GHz observation for our analyses.
@@
The mean flux of J1818 at each frequency is shown in the top panel of Figure 4. The light gray bars denote the frequency ranges of the 9- and 22- GHz datasets, to distinguish them from the other observations. We note that these observations are spread out over the course of roughly 100 days, and given the known flux variability of magnetars over time as well as the long-term emission mode changes reported by Rajwade et al. (2022) and Fisher et al. (2024), these results cannot be used to study the instantaneous broadband spectrum of J1818 nor the time variability of its spectrum.
@@ 3.3 Phase-resolved Spectroscopy
In order to better understand the emission properties of the individual profile components, we also measured the time-averaged spectral index of each component. We defined the on-pulse region of each individual component using the highlighted phase ranges in Figure 1, and calculated mean on-pulse fluxes for each component by using only these relevant parts of the profile. We used these on-pulse fluxes to calculate the spectral indices of the individual profile components at each frequency. The mean fluxes of each component are shown in the bottom panel of Figure 4, and the resulting fitted spectral indices are provided in Table 2.
@@ 4.3 Polarization
Figure 12: The distribution of linear polarization fraction (L/I) and circular polarization fraction (C/I) for each single pulse at 820 MHz. 1-$\sigma$ error bars on each point are denoted in light gray. The mean values of the polarization fractions are shown in the histograms with black dashed lines.
Figure 13: The distribution of linear polarization fraction (L/I) and circular polarization fraction (C/I) for each single pulse at 9 GHz. 1-$\sigma$ error bars on each point are denoted in light gray. The mean values of the polarization fractions are shown in the histograms with black dashed lines.
@@
We examined the degree of linear and circular polarization within each single pulse by measuring the polarized flux and comparing it with the flux of the total-intensity pulse. The errors on the linearly and circularly polarized fluxes are calculated as the off-pulse RMS of the respective polarized pulse profile (i.e. the RMS of the linearly polarized profile is the error on the linearly polarized flux). To examine the distribution of polarization fractions, we first excised pulses from our sample which had insufficient S/N to adequately measure the flux.
The resulting distributions of single pulse polarization fractions (linear, L/I, and circular, C/I) at 820 MHz and 9 GHz are respectively shown in Figures 12 and 13. The mean values ($\mu$) of the single pulse polarization fractions are denoted in the histograms. In comparison to the polarization fractions of the time-integrated profiles given in Table 2, the pulses at 820 MHz have a lower average L/I and higher average C/I. At 9 GHz, the L/I distribution is wider and there are more highly linearly polarized pulses, but the C/I distributions are similar at both frequencies.
At both frequencies, a portion of the single pulses have the sign of the circular polarization reversed. Pulse-to-pulse changes in the handedness of the circular polarization have been observed before in radio magnetars alongside structures in the PA which deviate significantly from the RVM (Kramer et al., 2007; Desvignes et al., 2024; Lower et al., 2024). Similar behavior was also observed in a long-period radio transient with similar emission properties to FRBs and magnetars (Men et al., 2025). We show some of the 9-GHz single pulses and their PA swings in Figure 14.
@@ 4.4 Searches for Periodic Substructure
For each of these pulses and their associated quasi-periodicities, we also wanted to measure a statistical significance for each quasi-periodicity. We took each pulse with an identified quasi-periodicity and randomly scrambled the order of its phase bins 50000 times. We took the ACF of each scrambled pulse and identified the maximum value outside of the first and last bin (e.g. the quasi-periodicity of the simulated pulse), and summed the values of the ACF at the maximum value and its harmonics. The simulated ACF sums form a null distribution, and the mean and standard deviation of this distribution can be compared to the summed ACF power of the real quasi-periodicity to obtain a significance ($\sigma$) for the detected quasi-periodicity.
Of the pulses selected for autocorrelation analysis, 108 pulses at 6 GHz and 15 pulses at 9 GHz had detected quasi-periodicities with significances above 4 $\sigma$. The distribution of quasi-periodicities at both frequencies is shown in Figure 16. At 6 GHz, the geometric mean $\mu_{g}$ of the detected quasi-periodicities is $4.9^{+2.0}_{-1.4}$ ms; at 9 GHz, the mean is $6.9^{+2.7}_{-1.9}$ ms, where the 1-$\sigma$ error bars are calculated using the geometric standard deviation $\sigma_{g}$; that is, the error bars range from $\mu_{g}/\sigma_{g}$ to $\mu_{g}\times\sigma_{g}$. The power-law scaling relation derived by Kramer et al. (2024) predicts a mean quasi-periodicity of $1.27\pm 0.08$ ms, while the distribution of their detected quasi-periodicities had a geometric mean of $2.37\pm 0.04$ ms.
The quasi-periodicities detected at 6 GHz by Kramer et al. (2024) were all shorter than 5 ms, at odds with the longer quasi-periodicities distributed over a larger range which we observe at both frequencies. In both radio pulsars and magnetars, a large spread of detected quasi-periodicity values have been measured (Mitra et al. 2015; Kramer et al. 2024). In particular, a clear bimodal distribution of detected periodicities was reported by Kramer et al. (2024) for the radio magnetar XTE J1810–197. We find that at both frequencies, the shorter quasi-periodicities have higher statistical significance than the longer quasi-periodicities. If the detection threshold were raised to significances $\geq 8$, the mean periodicities would become $3.8^{+1.4}_{-1.0}$ ms at 6 GHz (from 44 pulses), and $6.2^{+2.1}_{-1.6}$ ms at 9 GHz (from seven pulses). Our results show a clear temporal variation in the periodic substructure of magnetars over time– regardless of the statistical significance threshold used, the mean quasi-periodicity of our detected pulses is higher than those preferred by Kramer et al. (2024).
It is also worth noting that for all of the radio magnetars studied by Kramer et al. (2024), the geometric mean of their quasi-periodicities is larger than that predicted by the fitted power-law scaling relationship. While their periodicities still agreed with the scaling relationship within one to two geometric standard deviations, further study would be necessary to determine whether the periodic sub-structure relationship differs for magnetars as opposed to normal pulsars. The periodicity of the sub-structure appears to only depend on the underlying rotational period of the neutron star for a diverse range of neutron star types observed over a wide range of observing frequencies. While our results are an outlier compared to the predictions of the scaling relationship proposed by Kramer et al. (2024), a larger sample of pulses over a range of observing epochs would be necessary to draw any conclusions about the nature of quasi-periodic substructure in magnetar single pulses.
@@
Figure 16: The distribution of detected quasi-periodicities in the single pulses of J1818 at 6 GHz (top) and 9 GHz (bottom). The observed and predicted means of the quasi-periodicities found by Kramer et al. (2024) are denoted with dashed red and green lines, respectively. The geometric mean and associated error bars for our distribution of detected quasi-periodicities are denoted with black lines and shaded in orange.